OpenStudy (anonymous):

Whats the inverse f(x)= -2x-9 ?

sam (.sam.):

f(x)= -2x-9 let f(x)=y y=-2x-9 Solve for x in terms of y,

OpenStudy (anonymous):

x----> -2x-9 ? <---y y=-2x-9 x=y+9/(-2) therefore, the ans is -(y+9)/2

OpenStudy (anonymous):

(-x+9)/2 in terms of x

OpenStudy (anonymous):

Thanks guys !